A cross-country skier skis 1.00km north and then 2.00km east on a horizontal snowfield. how far and in what direction is she from the starting point?
1
Expert's answer
2019-06-26T08:58:05-0400


a) We can find the resultant displacement of the cross-country skier from the Pythagorean theorem:


"R = \sqrt{R_x^2 + R_y^2} = \sqrt{(2.0 km)^2 + (1.0 km)^2} = 2.24 km."

b) We can find the direction of the displacement vector from the geometry of the problem:


"tan\theta = \dfrac{R_x}{R_y},""\theta = tan^{-1} (\dfrac{R_x}{R_y}) = tan^{-1} (\dfrac{2.0 km}{1.0 km}) = 63.4^{\circ}."

Answer:

a) "R = 2.24 km."

b) "\theta = 63.4^{\circ}."